how to solve system by elimination for example 2x-10y=24
-4x+10y=-8

Answer by checkley77(12569)

2x-10y=24
-4x+10y=-8 add.
---------------
-2x=16
x=16/-2
x=-8 ans.
2*-8-10y=24
-16-10y=24
-10y=24+16
-10y=40
y=40/-10
y=-4 ans.
Proof:
-4*-8+10*-4=-8
32-40=-8
-8=-8